OpenStudy (anonymous):

Two planes, which are 3800 miles apart, fly toward each other. Their speeds differ by 50 mph. If they pass each other in 4 hours, what is the speed of each?

OpenStudy (anonymous):

Plane 1 is traveling at 450 MPH Plane 2 is traveling at 500 MPH

OpenStudy (anonymous):

4R + 4(R+50) = 3800

OpenStudy (anonymous):

Let R = the speed traveled by Plane 1.

OpenStudy (anonymous):

Apply the distance forumla: d=rt . Let x = faster plane then x-50 = slower plane . 4x + 4(x-50) = 3800 Dividing both sides by 3: 4x + 4x-200 = 3800 8x = 3800+200 x = 4.000/8 x = 500 mph (faster plane) . Slower plane: x-50 = 500-50 = 450 mph

OpenStudy (anonymous):

Thank ya'll